# Directed Acyclic Outerplanar Graphs Have Constant Stack Number

### v2

@@ 3.2 Directed $\boldsymbol{H}$-Partitions
The assumption in Lemma 3.8 that the edges $e_{1},\ldots,e_{k}$ can be covered with at most $p$ directed paths in $G[v]$ can be easily relaxed. Indeed if $B$ is the parent block of $v$ in the block-cut tree of $H$, it is enough that the $s$-stack layout of $G[v]$ as part of the $s$-stack layout of $G[B]$ in the statement of the Lemma contains a set of at most $m$ non-crossing matchings that cover $e_{1},\ldots,e_{k}$. Then the above proof can be easily adapted to show $\operatorname{sn}(G)\leq 2smt$. Having at most $p$ directed paths is clearly enough to have $m\leq 2p$ non-crossing matchings in every topological vertex ordering of $G[v]$.
@@ 3.3 Directed $\boldsymbol{H}$-Partitions of Acyclic Outerplanar Graphs
The goal of this section is to construct directed $H$-partitions $\mathcal{P}$ for every outerplanar DAG $G$, such that we can apply Lemma 3.8 from the previous section to reason that $\operatorname{sn}(G)$ is bounded by a constant. In particular, we aim for a block-monotone $H$, with each part $P\in\mathcal{P}$ inducing a relatively simple subgraph in $G$, as well as small cut cover numbers. Instead of bounding the cut cover number globally, it suffices to have it constant for each block of $H$ locally. Formally, if $B$ is a block of $H$, then $\mathcal{P}_{B}=\{P\in\mathcal{P}\mid v_{P}\in B\}$ is a directed $B$-partition of the corresponding subgraph $G[B]$ of $G$, and we want that the cut cover number of $\mathcal{P}_{B}$ is constant. This is enough to apply Lemma 3.10 from the previous section.
@@
    For each $P\in\mathcal{P}$ the graph $G[v_{P}]$ contains two directed paths $Q_{1},Q_{2}$ such that every vertex of $G-P$ that is stacked onto an edge in $G[v_{P}]$ is stacked onto an edge of $Q_{1}$ or $Q_{2}$.
@@
We divide the proof into five parts. First we define $\mathcal{P}$ according to property (P1). Second, we analyze the subgraph of $G$ induced by the vertices of each part $P\in\mathcal{P}$ and thereby verify property (P2). Third, we show that $\mathcal{P}$ is indeed a directed $H$-partition, i.e., the quotient $H:=G/\mathcal{P}$ is a well-defined directed graph. Then we show (P3), i.e., that $H$ is block-monotone. At last, we prove that the cut cover number for each block of $H$ is at most $4$, verifying property (P4).
@@
Construction of $\boldsymbol{\mathcal{P}}$. For each monotone vertex $u$ of $G$, let $P(u)$ be the set of vertices in the transitive subgraph below $u$. By definition we have $u\in P(u)$ and that $u$ is the only monotone vertex in $P(u)$. Moreover recall that the root of the construction tree $T$ is a monotone vertex and thus every transitive vertex of $G$ lies in $P(u)$ for some monotone $u$. Hence $\mathcal{P}=\{P(u)\mid u\text{ monotone vertex of $G$}\}$ is indeed a partition of $V(G)$ satisfying (P1).
For the remainder of the proof it will be convenient to consider a construction sequence of $G$ in which every monotone vertex $u$ is immediately followed by the vertices in the transitive subgraph below $u$. We consider such sequence vertex by vertex and argue about intermediate versions of $G$ and $\mathcal{P}$ (and thus of $H=G/\mathcal{P}$). At the beginning we have only the base edge $xy$ directed from $x$ to $y$, which are both labeled M. Then $\mathcal{P}=\{P(x),P(y)\}$ with $P(x)=\{x\}$ and $P(y)=\{y\}$. In each subsequent step, a vertex $u$ is stacked onto an edge $vw$. The parent of $u$ in the construction tree $T$ is the younger among $v,w$. If $u$ is a transitive vertex (i.e., labeled T in $T$), then $u$ is simply added to the part $P$ in $\mathcal{P}$ that contains the parent of $u$ in $T$. Otherwise, if $u$ is a monotone vertex (i.e., labeled M in $T$), then a new part $P(u)$ is added to $\mathcal{P}$ consisting of only $u$. Figure 11 shows an example of the resulting partition. Note that this iterative process indeed eventually results in the partition $\mathcal{P}=\{P(u)\mid u\text{ monotone vertex of $G$}\}$ of $G$ as described above.
Figure 11: Left: An outerplanar DAG with base edge $xy$ and an $H$-partition $\mathcal{P}$. The vertex $u$ is stacked onto the edge $vw$ and is the monotone vertex in $P(u)$. The paths $Q_{1}^{+}$ and $Q_{2}^{+}$ for $P(u)$ are drawn thick. The vertices marked with red circles certify that the cut cover number of the part $P(u)$ is at most 4. In a later step, $z$ is stacked onto $ab$, introducing a bridge that is a new block in $H$. Right: the quotient $H=G/\mathcal{P}$
@@
$\boldsymbol{\mathcal{P}}$ fulfills (P2). We shall argue that each part $P\in\mathcal{P}$ fulfills (P2) by showing this for the moment when $P$ is created in the construction sequence and maintaining (P2) for $P$ whenever $P$ is augmented with a new vertex thereafter. So fix a monotone vertex $u$ and consider the step in the construction sequence when the part $P=P(u)$ is created. As $P$ is created with just the single vertex $u$, property (P2) holds vacuously immediately after its creation. And if $u=x$, then $P=P(u)$ will never be augmented with a new vertex and (P2) holds throughout. To show that (P2) is maintained for $P=P(u)$ for $u\neq x$ with each step of the construction sequence, we maintain two directed paths $Q_{1}^{+},Q_{2}^{+}$ associated to $P$ and containing all edges induced by $P$ onto which a vertex can be stacked (see again Figure 11). If $u=y$, we initialize $Q^{+}_{1}:=(v,u)$ with $v=x$ and $Q^{+}_{2}:=(w,u)$ with $w=x$. If $u\neq x,y$ and the monotone vertex $u$ is stacked onto edge $vw$, we initialize $Q_{1}^{+}:=(v,u)$ and $Q_{2}^{+}:=(w,u)$. In any case, we have a $u$-$v$-path $Q_{1}^{+}$ and a $u$-$w$-path $Q_{2}^{+}$, and in $G$ both paths are either consistently oriented towards $u$ or both consistently oriented away from $u$. It holds that the next vertex in the construction sequence with a parent in $P$ is stacked onto an edge of $Q_{1}^{+}$ or an edge of $Q_{2}^{+}$. Moreover, the subpaths $Q_{1}:=Q_{1}^{+}-\{v\}$ and $Q_{2}:=Q_{2}^{+}-\{w\}$ are contained in $G[v_{P}]$ and trivially fulfill (P2) for $P=P(u)$.
Now consider the next step with a transitive vertex $z$ added to $P=P(u)$. (If the next vertex is monotone, then $P$ is final and we are done.) Then $z$ is stacked onto an edge $ab$ of $Q_{1}^{+}$ or $Q_{2}^{+}$. We replace edge $ab$ in $Q_{1}^{+}$ (or $Q_{2}^{+}$) by the path $(a,z,b)$. This way, $Q_{1}^{+}$ and $Q_{2}^{+}$ are still oriented in $G$ consistently towards $u$ or away from $u$. Since $G$ is outerplanar, no future vertex is also stacked onto $ab$. Hence, the next vertex with a parent in $P$ is again stacked onto an edge of $Q_{1}^{+}$ or an edge of $Q_{2}^{+}$. It follows that the subpaths $Q_{1}:=Q_{1}^{+}-\{v\}$ and $Q_{2}:=Q_{2}^{+}-\{w\}$ are contained in $G[v_{P}]$ and again fulfill (P2) for $P=P(u)$.
[…]
$\boldsymbol{\mathcal{P}}$ is a directed $\boldsymbol{H}$-partition. We show that in each step of the construction sequence $\mathcal{P}$ is indeed a directed $H$-partition, i.e., all edges between any two parts $P(u_{1}),P(u_{2})\in\mathcal{P}$ are oriented in the same direction. Assume without loss of generality that monotone vertex $u_{1}$ appears before monotone vertex $u_{2}$ in the construction sequence. Thus at the time $u_{2}$ is stacked onto some edge, $P(u_{1})$ is already in $\mathcal{P}$ and $u_{2}$ is a child of some vertex in $P(u_{1})$. If $u_{2}$ is a right (left) child, we show that all edges are oriented from $P(u_{1})$ to $P(u_{2})$, respectively from $P(u_{2})$ to $P(u_{1})$. This clearly holds immediately after the construction step for $u_{2}$.
By symmetry, assume that $u_{2}$ is a right child. Then the paths $Q_{1}^{+}$ and $Q_{2}^{+}$ for $P(u_{2})$ are both in $G$ directed towards $u_{2}$ and away from the parents of $u_{2}$. Now consider the next step with a transitive vertex $z$ added to $P(u_{2})$ such that $z$ also has a parent in $P(u_{1})$. Then $z$ is stacked onto the first edge of $Q_{1}^{+}$ or the first edge of $Q_{2}^{+}$. As $z$ is transitive, the edge between $P(u_{1})$ and $z$ is oriented towards $z\in P(u_{2})$, as desired.
[…]
$\boldsymbol{H}$ is a block-monotone DAG. To show that (P3) holds, we start by observing that $H=G/\mathcal{P}$ is a DAG. By the previous paragraph, $H$ is a well-defined directed graph, but we need to argue that it is acyclic999 The quotient of an outerplanar DAG obtained by contracting some edges might be cyclic. For example consider a $6$-cycle with alternating edge orientations. Contracting a maximum matching results in a directed $3$-cycle. . Since all edges between two parts of $\mathcal{P}$ are oriented consistently, it is sufficient to verify that $H$ remains acyclic whenever a new part $P$ is created. So $P=P(u)$ for some monotone vertex $u$ and either both edges incident to $u$ are oriented towards $u$ or both edges are oriented away from $u$. Thus, $v_{P}$ is not on any cycle in $H$.
Further observe that each part in $\mathcal{P}$ induces a connected subgraph of $G$. Thus, the quotient graph $H=G/\mathcal{P}$ is a minor of $G$. In particular, $H$ is outerplanar and every block of $H$ is maximal outerplanar.
It remains to show for (P3) that $H$ is block-monotone. Again we do this using the construction sequence of $G$. Initially, there is only the base edge $xy$ and we have $\mathcal{P}=\{P(x),P(y)\}$ with $P(x)=\{x\}$ and $P(y)=\{y\}$. We declare the currently only edge of $H$ to be the base edge for the currently only block of $H$.
[…]
Each block has cut cover number at most $\boldsymbol{4}$. Let $B$ be a block of $H$ and $G[B]\subseteq G$ be the corresponding subgraph of $G$. Further, let $P$ be a part of $\mathcal{P}_{B}$, i.e., such that $v_{P}$ lies in $B$. The goal is to show that part $P$ has cut cover number at most $4$, i.e., to find a set $S$ of at most four vertices in $G[B]$ that cover all edges of $G[B]$ with one endpoint in $P$ and the other endpoint in $B$ but in another part of $\mathcal{P}_{B}$. Recall that the endpoints of every edge in $G$ are in a parent/child relation. In particular, we shall consider edges $e\in E(G[B])$ whose parent-endpoint lies in $P$ while the child-endpoint does not, and the edges whose child-endpoint is in $P$ while the parent-endpoint is not.
Let $u$ be the monotone vertex of $G$ with $P=P(u)$. Clearly, if $u=x$, then $P(u)=\{u\}$ and it is enough to take $S=\{u\}$. If $u\neq x$, let $Q_{1}^{+}$ and $Q_{2}^{+}$ be the paths associated to $P(u)$. By symmetry assume that $Q_{1}^{+}$ is a directed $v$-to-$u$-path and $Q_{2}^{+}$ is a directed $w$-to-$u$-path in $G$ (either because $u=y$ or $u\neq y$ is some right child). Let us first assume that $G[B]$ contains $v$ and $w$. In this case let $S$ be the set consisting of $v$, $w$, the neighbor of $v$ in $Q_{1}^{+}$, and the neighbor of $w$ in $Q_{2}^{+}$, see again Figure 11. (Several of these vertices might coincide.) Then $|S|\leq 4$ and we claim that $S$ covers every edge in $G[B]$ with exactly one endpoint in $P$. Indeed, every edge in $G[B]$ with child-endpoint in $P$ but parent-endpoint outside $P$ has as parent-endpoint either $v$ or $w$. So let $e$ be an edge with parent-endpoint in $P$ but child-endpoint $z$ in $G[B]-P$. Hence the part $P^{\prime}$ in $\mathcal{P}$ containing $z$ was created later in the construction sequence than $P$. As $P^{\prime}$ also belongs to the block $B$, vertex $v_{P^{\prime}}$ has a second neighbor in $B$ at the time of its creation. Thus the corresponding monotone vertex was stacked onto an edge with one endpoint in $P$ and the other endpoint in another part of $\mathcal{P}_{B}$, and hence has a parent in $S$.
Finally assume that $G[B]$ does not contain $v$ and $w$ (e.g., if $B$ is the bridge in Figure 11). Then $v_{P}$ is a cut vertex of $H$ and $B$ is a child block of $v_{P}$ in the block-cut tree of $H$. Thus every edge in $G[B]$ with exactly one endpoint in $P$ has the parent-endpoint in $P$ and the child-endpoint in $G[B]-P$. Recall that block $B$ was initialized as a bridge when a monotone vertex $z$ was stacked onto an edge $ab$ of $Q_{1}$ or $Q_{2}$. In this case let $S=\{a,b\}$. Then $|S|=2$ and $S$ clearly contains the parent-endpoint of every edge in $G[B]$ with exactly one vertex in $P$.
3.4 A Constant Upper Bound
By (P1) of Lemma 3.12, each part of the constructed directed $H$-partition induces the transitive subgraph below a monotone vertex. Bounding the stack number of these subgraphs is the last missing piece before we prove Theorem 1.3.
Lemma 3.14.
Let $G$ be a maximal outerplanar DAG with a fixed base edge and $T$ be its (rooted) construction tree. Then the transitive subgraph below every monotone vertex $w$ admits a 1-stack layout.
Figure 12: An outerplanar DAG with its construction tree and a 1-stack layout of the transitive subgraph below $w$
Proof 3.15.
Let $uv$ denote the parent edge of $w$. By symmetry, we assume that $w$ is a right child. First recall that $w$ has at most two children $x$ and $y$ in $T$, corresponding to the children of $uw$, respectively $vw$, in $G$. Let $G_{x}$ ($G_{y}$) denote the subgraph of $G$ induced by $u$ ($v$), $w$, and the transitive subgraph below $x$ ($y$), see Figure 12. Note that the union of $G_{x}$ and $G_{y}$ contains the transitive subgraph below $w$. Now observe that $G_{x}$ and $G_{y}$ are transitive outerplanar DAGs, i.e., obtained from $uw$, respectively $vw$, by repeatedly stacking transitive children. Therefore, they have a unique topological ordering which coincides with the ordering of the vertices around the outer face, starting with the tail of the base edge and ending with its head. It is well known that one stack suffices for outerplanar graphs with this vertex ordering. A 1-stack layout for $G_{x}\cup G_{y}$ is now obtained by concatenating the layouts of $G_{x}-w$ and $G_{y}$, which in particular gives a 1-stack layout for the transitive subgraph below $w$.
Finally, we are ready to prove the first main result of this paper.
See 1.3

### v3

@@ 3.2 Directed $\boldsymbol{H}$-Partitions
The assumption in 3.9 that the edges $e_{1},\ldots,e_{k}$ can be covered with at most $p$ directed paths in $G[v]$ can be easily relaxed. Indeed, if $B$ is the parent block of $v$ in the block-cut tree of $H$, it is enough that the $s$-stack layout of $G[v]$ as part of the $s$-stack layout of $G[B]$ in the statement of the Lemma contains a set of at most $m$ non-crossing matchings that cover $e_{1},\ldots,e_{k}$. Then the above proof can be easily adapted to show $\operatorname{sn}(G)\leq 2smt$. Having at most $p$ directed paths is clearly enough to have $m\leq 2p$ non-crossing matchings in every topological vertex ordering of $G[v]$.
@@ 3.3 Directed $\boldsymbol{H}$-Partitions of Acyclic Outerplanar Graphs
The goal of this section is to construct directed $H$-partitions $\mathcal{P}$ for every outerplanar DAG $G$, such that we can apply Lemma 3.8 from the previous section to reason that $\operatorname{sn}(G)$ is bounded by a constant. In particular, we aim for a block-monotone $H$, with each part $P\in\mathcal{P}$ inducing a relatively simple subgraph in $G$, as well as small cut cover numbers. Instead of bounding the cut cover number globally, it suffices to have it constant for each block of $H$ locally. Formally, if $B$ is a block of $H$, then $\mathcal{P}_{B}=\{P\in\mathcal{P}\mid v_{P}\in B\}$ is a directed $B$-partition of the corresponding subgraph $G[B]$ of $G$, and we want that the cut cover number of $\mathcal{P}_{B}$ is constant. This is enough to apply 3.9 from the previous section. These properties of the following lemma are illustrated in Figure 6 and then combined in Figure 7.
@@
    For each $P\in\mathcal{P}$ the graph $G[P]$ contains two directed paths $Q_{1},Q_{2}$ such that every vertex of GPG-P that is stacked onto an edge in $G[P]$ is stacked onto an edge of $Q_{1}$ or $Q_{2}$.
@@
Figure 6: The properties guaranteed by Lemma 3.12. Figure 7 shows a larger example combing all properties. 1 Each monotone vertex (thick) has its own part (orange) including all transitive vertices until the next monotone vertex, i.e., the transitive subgraph below it (right). 2 Two paths $Q_{1}^{+},Q_{2}^{+}$ cover all outer edges of part $P$. Adding another transitive vertex $z$ extends the path (dashed). Removing $v$ and $w$ yields $Q_{1}$ and $Q_{2}$. 3 Each block of the quotient (right) consists of a base edge (thick) and monotone vertices corresponding to the monotone vertices in $G$ (left). 4 Inside each block, the cut cover number is bounded, i.e., for each part $P$ there are four vertices (circled red) that cover all edges leaving $P$.
@@
We divide the proof into five parts. First we define $\mathcal{P}$ according to property 1. Second, we analyze the subgraph of $G$ induced by the vertices of each part $P\in\mathcal{P}$ and thereby verify property 2. Third, we show that $\mathcal{P}$ is indeed a directed $H$-partition, i.e., the quotient $H:=G/\mathcal{P}$ is a well-defined directed graph. Then we show 3, i.e., that $H$ is block-monotone. At last, we prove that the cut cover number for each block of $H$ is at most $4$, verifying property 4.
{subproof}
[Construction of $\boldsymbol{\mathcal{P}}$] For each monotone vertex $u$ of $G$, let $P(u)$ be the set of vertices in the transitive subgraph below $u$. By definition, we have $u\in P(u)$ and that $u$ is the only monotone vertex in $P(u)$. Moreover recall that the root of the construction tree $T$ is a monotone vertex and thus every transitive vertex of $G$ lies in $P(u)$ for some monotone $u$. Hence $\mathcal{P}=\{P(u)\mid u\text{ monotone vertex of $G$}\}$ is indeed a partition of $V(G)$ satisfying 1.
For the remainder of the proof it will be convenient to consider a construction sequence of $G$ in which every monotone vertex $u$ is immediately followed by the vertices in the transitive subgraph below $u$. We consider such a sequence vertex by vertex and argue about intermediate versions of $G$ and $\mathcal{P}$ (and thus of $H=G/\mathcal{P}$). At the beginning we have only the base edge $xy$ directed from $x$ to $y$, which are both labeled M. Then $\mathcal{P}=\{P(x),P(y)\}$ with $P(x)=\{x\}$ and $P(y)=\{y\}$. In each subsequent step, a vertex $u$ is stacked onto an edge $vw$, where $v,w$ are the parents of $u$. Recall, however, that the parent of $u$ in the construction tree $T$ is the younger among v,wv,w. [^10] We remark that, unless stated otherwise, parent/child refers to the construction sequence. If $u$ is a transitive vertex (i.e., labeled T in $T$), then $u$ is simply added to the part $P$ in $\mathcal{P}$ that contains the parent of $u$ in $T$. Otherwise, if $u$ is a monotone vertex (i.e., labeled M in $T$), then a new part $P(u)$ is added to $\mathcal{P}$ consisting of only $u$. Figures 7 and 6 show an example of the resulting partition. Note that this iterative process indeed eventually results in the partition $\mathcal{P}=\{P(u)\mid u\text{ monotone vertex of $G$}\}$ of $G$ as described above.
Figure 7: Full example for the proof of Lemma 3.12. For better readability, each step is shown separately in Figure 6. Left: An outerplanar DAG with base edge $xy$ and a directed $H$-partition $\mathcal{P}$ (orange). The vertex $u$ is stacked onto the edge $vw$ and is the unique monotone vertex in $P(u)$. The paths $Q_{1}^{+}$ (darkblue from $v$ to $u$) and $Q_{2}^{+}$ (lightblue from $w$ to $u$) for $P(u)$ are drawn thick. The vertices marked with red circles certify that the cut cover number within the large block (all parts except for the rightmost) of the part $P(u)$ is at most 4. In a later step, $z^{\prime\prime}$ is stacked onto $ab$, introducing a bridge that is a new block in $H$. Right: The quotient $H=G/\mathcal{P}$, where each part of $\mathcal{P}$ is contracted to a single vertex. The vertices resulting from the parts $P(x),P(y),$ and $P(u)$ are labeled.
{subproof}
$\boldsymbol{\mathcal{P}}$ fulfills 2] We shall argue that each part $P\in\mathcal{P}$ fulfills 2 by showing this for the moment when $P$ is created in the construction sequence and maintaining 2 for $P$ whenever $P$ is augmented with a new vertex thereafter. So fix a monotone vertex $u$ and consider the step in the construction sequence when the part $P=P(u)$ is created. As $P$ is created with just the single vertex $u$, property 2 holds vacuously immediately after its creation. Moreover, if $u=x$, i.e., the tail of the base edge $xy$, then $P=P(u)$ will never be augmented with a new vertex and 2 holds throughout. To show that 2 is maintained for $P=P(u)$ for $u\neq x$ with each step of the construction sequence, we maintain two directed paths $Q_{1}^{+},Q_{2}^{+}$ associated to $P$ and containing all edges of the subgraph induced by $P$ onto which a vertex can be stacked (see again Figures 7 and 6). If $u=y$, we initialize both $Q^{+}_{1}$ and $Q^{+}_{2}$ with $(x,y)$. If $u\neq x,y$ and the monotone vertex $u$ is stacked onto edge $vw$, we initialize $Q_{1}^{+}:=(v,u)$ and $Q_{2}^{+}:=(w,u)$. In any case, we have a $u$-$v$-path $Q_{1}^{+}$ and a $u$-$w$-path $Q_{2}^{+}$ (taking $v=w=x$ if $u=y$), where in $G$ both paths are either consistently oriented towards $u$ or both consistently oriented away from $u$. It holds that the next vertex in the construction sequence with a parent in $P$ is stacked onto an edge of $Q_{1}^{+}$ or an edge of $Q_{2}^{+}$. Moreover, the subpaths $Q_{1}:=Q_{1}^{+}-\{v\}$ and $Q_{2}:=Q_{2}^{+}-\{w\}$ are contained in $G[P]$ and trivially fulfill 2 for $P=P(u)$.
Now consider the next step with a transitive vertex $z$ added to $P=P(u)$. (If the next vertex is monotone, then $P$ is final and we are done.) Then $z$ is stacked onto an edge $ab$ of $Q_{1}^{+}$ or $Q_{2}^{+}$ (gray/dashed in Figure 6). We replace edge $ab$ in $Q_{1}^{+}$ (or $Q_{2}^{+}$) by the path $(a,z,b)$. This way, $Q_{1}^{+}$ and $Q_{2}^{+}$ are still oriented in $G$ consistently towards $u$ or away from $u$. Since $G$ is outerplanar, no future vertex is also stacked onto $ab$. Hence, the next vertex with a parent in $P$ is again stacked onto an edge of $Q_{1}^{+}$ or an edge of $Q_{2}^{+}$. It follows that the subpaths $Q_{1}:=Q_{1}^{+}-\{v\}$ and $Q_{2}:=Q_{2}^{+}-\{w\}$ are contained in $G[P]$ and again fulfill 2 for $P=P(u)$.
{subproof}
$\boldsymbol{\mathcal{P}}$ is a directed $\boldsymbol{H}$-partition] We show that in each step of the construction sequence $\mathcal{P}$ is indeed a directed $H$-partition, i.e., all edges between any two parts $P(u_{1}),P(u_{2})\in\mathcal{P}$ are oriented in the same direction. Assume without loss of generality that monotone vertex $u_{1}$ appears before monotone vertex $u_{2}$ in the construction sequence. Thus at the time $u_{2}$ is stacked onto some edge, $P(u_{1})$ is already in $\mathcal{P}$ and contains at least one parent of $u_{2}$. If $u_{2}$ is a right (left) child, we show that all edges are oriented from $P(u_{1})$ to $P(u_{2})$, respectively from $P(u_{2})$ to $P(u_{1})$. This clearly holds immediately after the construction step for $u_{2}$.
By symmetry, assume that $u_{2}$ is a right child. Then the paths $Q_{1}^{+}$ and $Q_{2}^{+}$ for $P(u_{2})$ are both in $G$ directed towards $u_{2}$ and away from the parents of $u_{2}$. Now consider the next step with a transitive vertex $z$ added to $P(u_{2})$ such that $z$ also has one of its parents in $P(u_{1})$. Then $z$ is stacked onto the first edge of $Q_{1}^{+}$ or the first edge of $Q_{2}^{+}$. As $z$ is transitive, the edge between $P(u_{1})$ and $z$ is oriented towards $z\in P(u_{2})$, as desired.
{subproof}
$\boldsymbol{H}$ is a block-monotone DAG] To show that 3 holds, we start by observing that $H=G/\mathcal{P}$ is a DAG. By the previous paragraph, $H$ is a well-defined directed graph, but we need to argue that it is acyclic [^11] The quotient of an outerplanar DAG obtained by contracting some edges might be cyclic. For example consider a $6$-cycle with alternating edge orientations. Contracting a maximum matching results in a directed $3$-cycle. . Since all edges between two parts of $\mathcal{P}$ are oriented consistently, it is sufficient to verify that $H$ remains acyclic whenever a new part $P$ is created. So we have $P=P(u)$ for some monotone vertex $u$ and either both edges incident to $u$ are oriented towards $u$ or both edges are oriented away from $u$. Thus, $v_{P}$ is not on any cycle in $H$.
Further observe that each part in $\mathcal{P}$ induces a connected subgraph of $G$. Thus, the quotient graph $H=G/\mathcal{P}$ is a minor of $G$. In particular, $H$ is outerplanar and every block of $H$ is maximal outerplanar. We refer to Figures 6 and 7 for examples of quotients.
It remains to show for 3 that $H$ is block-monotone. Again we do this using the construction sequence of $G$. Initially, there is only the base edge $xy$, and we have $\mathcal{P}=\{P(x),P(y)\}$ with $P(x)=\{x\}$ and $P(y)=\{y\}$. We declare the currently only edge of $H$ to be the base edge for the currently only block of $H$.
{subproof}
[Each block has cut cover number at most 4] Let $B$ be a block of $H$ and $G[B]\subseteq G$ be the corresponding subgraph of $G$. Further, let $P$ be a part of $\mathcal{P}_{B}$, i.e., such that $v_{P}$ lies in $B$. The goal is to show that part $P$ has cut cover number at most $4$ within the block $B$, i.e., to find a set $S$ of at most four vertices in $G[B]$ that cover all edges of $G[B]$ with one endpoint in $P$ and the other endpoint in $B$ but in another part of $\mathcal{P}_{B}$. Recall that the endpoints of every edge in $G$ are in a parent/child relation by construction of the outerplanar graph, where each child has two parents independent of the edge directions. In particular, we shall consider edges $e\in E(G[B])$ whose parent-endpoint lies in $P$ while the child-endpoint does not, and edges whose child-endpoint is in $P$ while the parent-endpoint is not.
(a) All edges with the child-endpoint in $P$ but the parent-endpoint not are incident to $v$ or $w$.
(b) All edges with the parent-endpoint in $P$ but the child-endpoint not are incident to $q_{1}$ or $q_{2}$. Note that stacking onto any edge with both endpoints in $P$ yields a new block and thus is not considered here, while stacking onto an edge with both endpoints not in $P$ creates only edges that do not intersect $P$.
(c) Left: If $G[B]$ does not contain $v,w$, we choose $S=\{a,b\}$. Again, recall that stacking onto any edge in part $P$ opens a new block. Right: The quotient with block $B$
Figure 8: Bounding the cut cover number in the proof of Lemma 3.12. Each subfigure refines Figure 7 to show the details in the different cases of the proof. Recall that the cut cover number is bounded for each block separately.
Let $u$ be the monotone vertex of $G$ with $P=P(u)$. Clearly, if $u=x$, then $P(u)=\{u\}$ and it is enough to take $S=\{u\}$. If $u\neq x$, let $Q_{1}^{+}$ and $Q_{2}^{+}$ be the paths associated to $P(u)$ as constructed above. By symmetry assume that $Q_{1}^{+}$ is a directed $v$-to-$u$-path and $Q_{2}^{+}$ is a directed $w$-to-$u$-path in $G$ (either because $u=y$ or $u\neq y$ is some right child). Let us first assume that $G[B]$ contains $v$ and $w$. In this case let $S$ be the set consisting of $v$, $w$, the neighbor $q_{1}$ of $v$ in $Q_{1}^{+}$, and the neighbor $q_{2}$ of $w$ in $Q_{2}^{+}$, see Figures 7 and 8. (Several of these vertices might coincide.) Then $|S|\leq 4$ and we claim that $S$ covers every edge in $G[B]$ with exactly one endpoint in $P$. Indeed, every edge in $G[B]$ with child-endpoint in $P$ but parent-endpoint outside $P$ has as parent-endpoint either $v$ or $w$ (Figure 8(a)). So let $e$ be an edge with parent-endpoint $p$ in $P$ but child-endpoint $z^{\prime}$ in G[B]PG[B]-P, see Figure 8(b). Hence the part $P^{\prime}$ in $\mathcal{P}_{B}$ containing $z^{\prime}$ was created later in the construction sequence than $P$. To show that $p$ is either $q_{1}$ or $q_{2}$ and therefore in $S$, we make two observations: First, recall that every monotone vertex that is stacked onto an edge with both endpoints in $P$ creates a new block. And second, observe that immediately after finishing $P$ in the construction sequence, the only two edges with exactly one endpoint in $P$ are $vq_{1}$ and $wq_{2}$, i.e., the first edge in each of $Q_{1}^{+}$ and $Q_{2}^{+}$. Together, it follows that whenever a new monotone vertex is introduced in the construction sequence, either it has both parents in $P$ and creates a new block, or it has exactly one parent in $P$, namely $q_{1}$ or $q_{2}$, or it has no parent in $P$. Since we only consider the block $B$, we conclude that all vertices introduced after finishing $P$ that have at least one parent in $P$ actually have exactly one parent in $P$, namely $q_{1}$ or $q_{2}$.
Finally, assume that $G[B]$ does not contain $v$ and $w$ (e.g., Figure 8(c) or if $B$ is the bridge in Figure 7). Then $v_{P}$ is a cut vertex of $H$ and $B$ is a child block of $v_{P}$ in the block-cut tree of $H$. In particular, $P$ is the first part in $G[B]$ according to the construction sequence, and thus every edge in $G[B]$ with exactly one endpoint in $P$ has the parent-endpoint in $P$ and the child-endpoint in G[B]PG[B]-P. Recall that block $B$ was initialized as a bridge when a monotone vertex $z^{\prime\prime}$ was stacked onto an edge $ab$ of $Q_{1}$ or $Q_{2}$. In this case let $S=\{a,b\}$. Then $|S|=2$ and $S$ clearly contains the parent-endpoint of every edge in $G[B]$ with exactly one vertex in $P$. Since we only need to bound the cut cover number of the partition restricted to some block $B$, this finishes 4.
We conclude that $\mathcal{P}$ is indeed a directed $H$-partition satisfying all four properties, which concludes the proof.
3.4 A Constant Upper Bound
By Property 1 of Lemma 3.12, each part of the constructed directed $H$-partition induces the transitive subgraph below a monotone vertex. Bounding the stack number of these subgraphs is the last missing piece before we prove Section 1.
Lemma 3.14.
Let $G$ be a maximal outerplanar DAG with a fixed base edge and let $T$ be its (rooted) construction tree. Then the transitive subgraph below every monotone vertex $w$ admits a 1-stack layout.
{sidefigure}[Uncaptioned image]
An outerplanar DAG with its construction tree and a 1-stack layout of the transitive subgraph below $w$
Proof 3.15.
Let $uv$ denote the parent edge of $w$. By symmetry, we assume that $w$ is a right child. First recall that $w$ has at most two children $a$ and $b$ in $T$, corresponding to the children of $uw$, respectively $vw$, in $G$. Let $G_{a}$ ($G_{b}$) denote the subgraph of $G$ induced by $u$ ($v$), $w$, and the transitive subgraph below $a$ ($b$), see Lemma 3.14. Note that the union of $G_{a}$ and $G_{b}$ contains the transitive subgraph below $w$. Now observe that $G_{a}$ and $G_{b}$ are transitive outerplanar DAGs, i.e., obtained from $uw$, respectively $vw$, by repeatedly stacking transitive children. Therefore, they have a unique topological ordering which coincides with the ordering of the vertices around the outer face, starting with the tail of their base edge and ending with its head. It is well known that one stack suffices for outerplanar graphs with this vertex ordering. A 1-stack layout for $G_{a}\cup G_{b}$ is now obtained by concatenating the layouts of $G_{a}-w$ and $G_{b}$, which in particular gives a 1-stack layout for the transitive subgraph below $w$.
Finally, we are ready to prove the first main result of this paper. This includes verifying the premises of 3.9, which we restate here for convenience.